The length of a rectangular sign is 4 feet longer than the width. If the sign's area is 480 square feet, find its length and width.
Thank you!!!!!!

Answer by rothauserc(4718)

let l be the length and w be the width
:
l = w + 4
:
l * w = 480
:
(w+4) * w = 480
:
w^2 +4w -480 = 0
:
(w+24) * (w-20) = 0
:
w = -24 and w = 20
:
we reject the negative value for w
:
************************************
length = 24 feet and width = 20 feet
************************************
: